# Can LLMs Understand Social Norms in Autonomous Driving Games?

### v1

@@ IV-B Scenario 2: Highway Platoon
1 10 10 1
2 10 10 1
3 10 10 1
4 10 10 1
TABLE IV: Highway Platoon Results
Figure 10: Number of Lane Changes
Figure 11: Time Step of Lane Changes
Figure 12: Time of Platoon
Figure 11 is the plot of the step time when the agents choose to change the lane, and figure 12 is the percentage of time of platoon in each game. From the plots, we observe that in most cases, agents opt to change lanes before time step 5. Moreover, in the majority of simulations, there is more than a 60 percent likelihood of two agents forming a platoon on the highway. From this, we can infer that LLM agents are inclined to change lanes and form platoons at an early stage, demonstrating an awareness that platoon formation is an efficient driving strategy on highways.
The behavior of changing lanes and forming platoons early is believed to mimic the social norms found in human society. Additionally, the varying behaviors of the agents may represent the diverse driving habits encountered in the real world. [argument of social norm in conclusion part?]
@@ V Conclusion And Future Work
This paper has demonstrated the potential of large language models (LLMs) in understanding and modeling social norms in autonomous driving scenarios. By introducing LLM-based agents into autonomous driving games, we have observed the emergence of social norms among agents navigating complex driving environments. Our experiments have shown that LLM-based agents can adapt and conform to social norms, thereby contributing to safer and more efficient driving behaviors. Overall, the advantage of employing LLM-based agents in autonomous driving games lies in their strong operability and analyzability, which facilitate experimental design and provide valuable insights into the dynamics of social norms in driving environments. By further investigating the behavior of LLM-based agents in various driving scenarios and refining our experimental framework, we can contribute to the development of more socially aware autonomous driving systems.
This work can be extended in the following ways: (1) We will utilize LLM-based agents in more complex real-world scenarios, such as sequential social dilemmas. (2) We aim to explore how to construct a unified, controllable, and efficient framework for simulating strategic interactions and facilitating game design using LLMs. (3) We will compare the behaviors of human players and LLM-based agents to determine the extent to which LLMs can achieve strategic reasoning.

### v2

@@ IV-B Scenario 2-Highway Platoon
1 50 46 0.92
2 50 49 0.98
3 50 43 0.86
4 50 40 0.80
TABLE III: Performance - highway
Table III summarizes the performance of LLM agents when varying numbers of background vehicles. The results show that the success rate of forming platoons is high, indicating the emergence of social norms among LLM agents. Figure 8 demonstrates lane-change behaviors. The histogram visualizes the frequency of lane-change actions. The x-axis denotes the number of lane-change actions and the y-axis is the number of simulations. We can tell that there is at most three lane-changes per simulation, and one lane-change in most cases. The green car changes lane more frequently in a comparison to the red car. Figure 9 is the stacked histogram of the time step when the agents choose to change the lane. It is shown that most of the LLM agents change lanes before the fourth time step. It indicates that LLM agents are inclined to change lanes and form platoons at an early stage, demonstrating an awareness that platoon formation is an efficient driving strategy on highways. Figure 10 plots the histogram of percentage time of platoon in each game. Moreover, in the majority of simulations where platoon is formed, there are more than a 60 percent of the time that two agents forming a platoon on the highway.
Figure 8: Lane-change frequency
Figure 9: Time step of lane-change
Figure 10: Time length of forming a platoon
@@ V Conclusion And Future Work
This paper has demonstrated the potential of LLMs in understanding and modeling social norms in autonomous driving scenarios. By introducing LLM agents into autonomous driving games, we have observed the emergence of social norms among agents navigating complex driving environments. Our experiments have shown that LLM agents can adapt and conform to social norms, thereby contributing to safer and more efficient driving behaviors. Overall, the advantage of employing LLM agents in autonomous driving games lies in their strong operability and analyzability, which facilitate experimental design and provide valuable insights into the dynamics of social norms in driving environments. By further investigating the behavior of LLM agents in various driving scenarios and refining our experimental framework, we can contribute to the development of more socially aware autonomous driving systems.
This work can be extended in the following ways: (1) We will utilize LLM agents in more complex real-world scenarios, such as sequential social dilemmas. (2) We aim to explore how to construct a unified, controllable, and efficient framework for simulating strategic interactions and facilitating game design using LLMs. (3) We will compare the behaviors of human players and LLM agents to determine the extent to which LLMs can achieve strategic reasoning.